Can you please provide the answers/tutorials to the mini-case for Chapter 17?

Mini-case:

a. What is a multinational corporation? Why do firms expand into other countries?

b. What are the six major factors that distinguish multinational financial management from financial management as practiced by a purely domestic firm?

c. Consider the following illustrative exchange rate.

US Dollars required to buy one unit of foreign currency - 1.2500 Euro

Units of foreign currency required to buy one US dollar - swedish krona - 7.0000

1. What is a direct quotation? what is the direct quote for euros?

2. What is an indirect quotation? What is the indirect quotation for kronor (the plural of krona is kronor)?

3. The euro and British pound usually are quoted as direct quotes. Most other currencies are quoted as indirect quotes. How would you calculate the indirect quote for a euro? How would you calculate the direct quote for a krona?

4. What is a cross rate? Calculate the two cross rates between euros and kronor.

5. Assume Possum Products can produce a package of jerky and ship it to France for $1.75. If the firm wants a 50% markup on the product, what should the jerky sell for in France?

6. Now assume that Possum Products begins producing the same package of jerky in France. The product costs 2 euros to produce and ship to Sweden, where it can be sold for 20 kronor. What is the dollar profit on the sale?

7. What is exchange rate risk?

d. Briefly describe the current international monetary system. How does the current system differ from teh system that was in place prior to August 1971?

e. What is a convertible currency? What problems arise when a multinational company operates in a country whose currency is not convertible?

f. What is the difference between spot rates and forward rates? When is the forward rate at a premium to the spot rate? At a discount?

g. What is interest rate parity? Currently, you can exchange 1 euro for 1.25 dollars in the 180-day forward market, and the risk-free rate on 180-day securities is 6% in the United States and 4% in France. Does interest rate parity hold? If not, which securities offer the highest expected return?

h. What is purchasing power parity? If a package of jerky costs $2 in the United States and purchasing power parity holds, what should be the price of the jerky package in France?

i. What effect does relative inflation have on interest rates and exchange rates?

j. Briefly discuss the international capital markets?

k. To what extent do average capital structures vary across different countries?

l. Briefly describe special problems that occur in multinational capital budgeting, and describe the process for evaluating a foreign project. Now consider the following project: A US company has the opportunity to lease a manufacturing facility in Japan for 2 years. The company must spend 1 billion initially to refurbish the plant. The expected net cash flows from the plant for the next 2 years, in millions, are CF1=Y500 and CF2=Y800. A similar project in the United States would have a risk-adjusted cost of capital of 10%. In the United States a 1-year government bond pays 2% interest and a 2-year bond pays 2.8%. In Japan, a 1-year bond pays 0.05% and a 2-year bond pays 0.26%. What is the project's NPV?

m. Briefly discuss special factors associated with the following areas of multinational working capital management.

1. Cash Management

2. Credit Management

3. Inventory Management
Answer #1

a.1. A Multinationational corporation is a firm that has its operations in number of countries (more than 1).

a.2. Firms expand to other countries :-

- to expand their markets where they could sell more for eg - Uber has expanded its operations from USA to all over the globe

- to have access to latest technology developed in foreign markets for eg - Many firms from developing countries have expanded to USA in order to avail latest technology

- to procure raw materials for eg - many Italian fashion firms have outsourced manufacturing facilities in China

- to diversify their risks so that it does not depend on a single market for its revenue

b. six major factors that distinguish multinational financial management from financial management as practiced by a purely domestic firm are:-

1. Foreign exchange eposure and risk- Domestic firms use limited financial instruments while multinational firms need to use financial instrument in order to manage their forex exposure

2. Corporate governance - Regulations and institutions are well known for domestic firms while for multinational firms, foreign countries regulation and institutional practices are uniquely different.

3. Culture, history and institutions - For domestic firms, each country has a known base case while for multinational firms, each foreign country is unique and not understood by MNE management

4. Political risk - There is negligible political risk for domestic firm while a multinational firm faces foreign exchange risks due to its subsidiaries, as well as import/export and foreign competitors.

5. Modification of domestic finance theories - Traditional finance theories apply for a domestic firm while Multinational firms must modify finance theories like capital budgeting and cost of capital because of foreign complexities

6. Modification of domestic finance instruments - there is limited use of financial instruments and derivatives because of fewer foreign exchange and political risk to a domestic firm while Multinational firms use modified financial instruments such as options, swaps, futures and letter of credit.

c.1.1 From a U.S. perspective, the quotes in the first column are called direct quotes because they are number of units of a foreign currency that can be purchased by 1 unit of the home currency.

c.1. 2. The direct quote for the euro is EUR/USD = 1.25. The financial press would report this as “the euro is trading at $1.25.”

c.2.1 Indirect quotations are the number of units of foreign currency that can be purchased with one unit of home currency (the home currency is the U. S. Dollar in this example).

c.2.2 The indirect quote for the krona is USD/SEK = 7.

c.3.1 Indirect quotations are the reciprocal of the direct quotation, and direct quotations are the reciprocal of the indirect quotation.

c.3.2 The indirect quote for the euro is USD/EUR = 1/1.25 = 0.80. The direct quote for the krona is SEK/USD = 1/7 = 0.1429.

c.4.1 The exchange rate between any two currencies which does not involve U. S. Dollars is a cross rate.

c.4.2 Use the exchange rates versus the dollar, so use the direct for the dollars per euro and the indirect for kronor per dollar. Here is the cross rate for kronor per euro: Cross rate = (Kronor / Dollar) X (Dollars / Euros ) = 7 × 1.2500 = 8.750 kronor per euro. Euros per krona cross rate = 1/(8.750 kronor per euro) = 0.1143 euros per krona

c.5.1 To achieve the markup, the price in dollars must be ($1.75)(1.50) = $2.625. The reported quote is the direct quote for dollars per euro. Therefore, the calculation using the direct quote is: Price in euros = 2.625 dollars / (1.25 dollars /euro) = (2.625 / 1.25 ) X (dollars /1 ) X (euros / dollars ) = 2.10 euros

c.6.1 Using the unrounded cross rate of 8.50 kronor per euro, we get: (2.0 euros)(8.50 kronor/euro) = 17.50 kronor The profit on the sale in Sweden is 20 – 17.50 = 2.50 kronor. Now, there are 0.1481 dollar per krona, so The dollar profit is (2.5 kronor)/ (7 kronor per dollar) = $0.36

c.7.1 The volatility inherent in a floating exchange rate system increases the uncertainty of cash flows that must be translated from one currency into another. This increase in uncertainty is exchange rate risk.

d.1 Prior to 1971, the world operated on a fixed exchange rate system. The value of the U. S. Dollar was linked to gold at the fixed price of $35 per ounce, and the values of other currencies were then tied to the dollar. For example, in 1964, the British pound was fixed at $2.80 for 1 pound, with a 1 percent permissible fluctuation around this rate. Thus, the British government had to regularly intervene in the foreign exchange market to keep the pound in the range of $2.77 to $2.83. When the pound fell, the Bank of England had to buy pounds, offering either foreign currencies or gold in exchange. Conversely, if the pound reached the top of the range, the Bank of England would sell pounds. The official exchange rates were occasionally “reset” to reflect changing economic conditions. The current international monetary system for most industrialized nations is a floating rate system. In this system, currency exchange rates are allowed to fluctuate in response to market conditions with a minimum of governmental intervention. Changes in currency demand can be due to trade deficits (i.e., one nation imports more from another nation than it exports, causing there to be higher relative demand for the currency of the bigger exporter). It can also be due to capital movements. For example, if interest rates are relatively high in one country, then investors might seek to purchase that country’s securities, which increases demand for that country’s currency. Central banks, like the U. S. Federal Reserve and the Bank of England, do intervene in the currency markets to smooth out fluctuations, but it is impossible for a central bank to permanently prop up a weak currency. Also, governments do enter into agreements to try to keep currencies within predetermined ranges. However, if market forces move the exchange rate outside one of these ranges, there is little that the countries can do other than adjust the target range. Sixteen participating countries in the European Monetary Union (as of Spring 2009) use the “euro.” The European Central Bank controls the monetary policy of the EMU nations using the Euro. Many countries still used a fixed exchange rate that is “pegged,” or fixed, with respect to another currency. Examples of pegged currencies are the Chinese yuan, which is pegged to the dollar and the Chad CFA franc, which is pegged to the French franc which is pegged to the euro. When a currency increases in value relative to another currency, it is said to appreciate. Under the fixed exchange rate system, strong currencies had to be revalued occasionally, which changed the tie to other currencies to a new, higher rate. Conversely, a currency that loses value is said to depreciate, and such currencies had to be devalued under the old fixed rate system.